OpenStudy (madmerc):

Which of the following is a solution of x^2 + 2x + 4? −2 + two i times the square root of 3 2 + two i times the square root of 3 1 − i times the square root of 3 −1 + i times the square root of 3

OpenStudy (madmerc):

@Nnesha can you help?

Nnesha (nnesha):

what's the quadratic formula ?

OpenStudy (madmerc):

y=mx+b @Nnesha

Nnesha (nnesha):

no that's the linear equation i'm talking about `quadratic formula `

Nnesha (nnesha):

x^2+2x+4 is a parabola not linear equation the highest degree `^2` makes it quadratic equation

OpenStudy (madmerc):

so what do I do? @Nnesha

Nnesha (nnesha):

\[\large\rm x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] quadratic formula should be in ur notes where A =leading coefficient b=coefficient of x term c=constant term plugin a , b ,c values \[\rm Ax^2+Bx+C=0\]

OpenStudy (madmerc):

so 2 + 2i- sqrt 3 @Nnesha

Nnesha (nnesha):

how did you get that show ur work what are a b c values ?

OpenStudy (madmerc):

2 and 4 are b and c @Nnesha

Nnesha (nnesha):

what did you plugin for a ?? :=))

OpenStudy (madmerc):

2 @Nnesha

Nnesha (nnesha):

no the equation is 1x^2+2x+4 what's a ? and post your work how did you get that answer

OpenStudy (madmerc):

i just guessed @Nnesha

Nnesha (nnesha):

don't guess it substitute a and b c for their values first tell me a= ? b=? c=?

OpenStudy (madmerc):

a is 1 b is 2 c is 4 @Nnesha

Nnesha (nnesha):

substitute a for 1 b for 2 andc for 4

OpenStudy (madmerc):

oh I was right! @Nnesha